Starting at home, Luis traveled uphill to the hardware store for 30 minutes at just 10 mph. He then traveled back home along the same path downhill at a speed of 30 mph.
What is his average speed for the entire trip from home to the hardware store and back?

Answer by Alan3354(69443)

RT avg = 2*10*30/(10+30) = 600/40
= 15 mi/hr